# SMART: Simulated Students Aligned with Item Response Theory for Question Difficulty Prediction

### v1

@@ 5.3 Baselines
Following prior work AlKhuzaey et al. (2024); Benedetto (2023); Benedetto et al. (2023); Kapoor et al. (2025a); North and Ormerod (2025), we adopt a strong finetuning baseline, Diff SFT, which finetunes a recent and capable encoder-based LLM, specifically ModernBERT-base Warner et al. (2024). We use a regression setup, minimizing the mean squared error loss, to directly predict item difficulties given the item passage and question. We also adopt a random forest baseline, which is common for difficulty prediction AlKhuzaey et al. (2024); Gombert et al. (2024), SBERT RF, where we encode the item passage and question text using SBERT Reimers and Gurevych (2019), specifically multi-qa-mpnet-base-dot-v1, and use the resulting embeddings as input features to fit a random forest regressor to item difficulty.
However, since our set of items is small, only containing 29 training samples per fold, we introduce several baselines that do not require training. We use a simple yet strong baseline, SBERT kNN Mean, that retrieves the $k$ items in the train set with the highest SBERT embedding cosine similarity, and takes their mean difficulty as the predicted test item difficulty. Similar methods have been used in automated scoring Bin et al. (2008); Gomaa and Fahmy (2012). We experiment with $k\in[1,5]$ on the validation set and find that $k=1$ works best, i.e., simply copying the difficulty of the most similar item. Finally, we adopt two LLM prompting baselines leveraging in-context learning (ICL) Brown et al. (2020), ICL Random and ICL kNN, where we prompt Llama-3.1-70B-Instruct with $k$ in-context examples of items retrieved from the train set at random or using k-Nearest Neighbors (kNN) Feng et al. (2024), respectively. We ask the model to first reason about the question’s difficulty in a chain-of-thought manner before generating a numeric difficulty prediction. For kNN retrieval, we use the same SBERT model as above, and find that $k=4$ works best out of $k\in[1,5]$ on the validation set.
@@
We compare SMART to other student simulation-based approaches for item difficulty prediction. For a fair comparison, we use the same base model and prompt template as SMART for the following methods, allowing these baselines to also serve as an ablation. First, we employ Response Gen Zero-shot, where we prompt an LLM in a zero-shot manner to generate a student response given the item and ability. Since the model has not been trained on the distribution of numeric student abilities, we convert abilities to corresponding labels from 10 descriptive buckets shown in Appendix A.2, ranging from “minimal” to “mastery”, with descriptions obtained by prompting Claude 3.5 Sonnet Anthropic (2024). This method is similar to the persona-based prompting method in He-Yueya et al. (2024), where we use ability instead of student demographics. Second, we employ Response Gen SFT, as explained in Section 4.1, which finetunes the LLM on ground-truth student responses given the item and ability. Prior works that finetune simulated students generally include prior student responses in the prompt He-Yueya et al. (2024); Zelikman et al. (2023); we use the student’s ability as a proxy for this information.
@@ 5.4 Experimental Setup
Model Item Difficulty Prediction Student Simulation
PCC $\uparrow$ SCC $\uparrow$ RMSE $\downarrow$ MAUVE $\uparrow$ FID $\downarrow$ Diversity KL $\downarrow$ $\theta$-Align $\uparrow$
Direct Item Difficulty Prediction Methods
Diff SFT (ModernBERT) $0.4937$ $0.3188$ $0.7627$
SBERT RF $0.4021$ $0.3115$ $0.8623$
SBERT kNN Mean $0.5959$ $0.5219$ $0.6806$
ICL Random (Llama-3.1-70B) $0.1215$ $0.0348$ $1.0426$
ICL kNN (Llama-3.1-70B) $0.5273$ $0.4717$ $0.7634$
Response Gen Zero-shot $0.2651$ $0.2436$ $0.9088$ $0.0058$ $0.2713$ $0.2535$ $0.0399$
Response Gen SFT $0.5621$ $0.4254$ $0.7174$ $0.1483$ $0.0847$ $\underline{0.0215}$ $0.4481$
SMART (ours) $\underline{0.6490}$ $0.4861$ $\underline{0.6699}$ $0.1315$ $0.0918$ $0.0482$ $\underline{0.7338}$
Response Gen Zero-shot $0.3446$ $0.3963$ $0.8568$ $0.0052$ $0.3386$ $0.4032$ $0.0211$
Response Gen SFT $0.6069$ 0.5830 $0.6905$ 0.1858 0.0757 0.0137 $0.5158$
SMART (ours) 0.6737 $\underline{0.5661}$ 0.6197 $\underline{0.1713}$ $\underline{0.0785}$ $0.0279$ 0.7655
Table 1: Cross-validation performance. Best performance is in bold and second best is underlined. SMART outperforms all other methods on difficulty prediction and shows high alignment with ability.
@@ 6 Results, Analysis, and Discussion
We now detail our quantitative results on student response simulation and item difficulty prediction, perform an ablation study, and qualitatively analyze the simulated student responses.
@@ 6.1 Quantitative Results
We report the average performance across all test items and across $5$ cross-validation folds in Table 1. Our approach, SMART, outperforms all other methods on all metrics, with the exception of SFT on SCC due to the variability of this metric on small test sets. Even with a small base LLM with 1B parameters, SMART outperforms all other methods on PCC and RMSE. ICL Random, our baseline which serves as a lower bound, has poor performance, indicating that even powerful LLMs are not able to perform this task without relevant context. On the other hand, ICL kNN and SBERT kNN Mean, by retrieving training items with high similarity with the target item, show improved performance over other direct prediction baselines. These methods likely leverage the fact that certain item types are more difficult than others, and we find that SBERT embeddings tend to have high similarity within item types.
@@
Diff SFT, which follows the approach of existing state-of-the-art methods AlKhuzaey et al. (2024); Benedetto (2023); Benedetto et al. (2023) to finetune an encoder LLM directly on IRT’s estimated item difficulty parameter, performs poorly on our data. This result is perhaps not surprising, since its training signal completely comes from given difficulty parameter values, and thus can only work when a large number of training items are available. In our low resource setting with few items to train with, student simulation is key: student ability and item difficulty are only related via the response score when training an IRT model, but our student simulator can effectively couple IRT training with textual, short-answer student responses and a scoring model. As a result, this extra information helps us to calibrate the IRT model better, making it easier to generalize to previously unseen items.
@@
We compare the performance of Response Gen Zero-shot, Response Gen SFT, and SMART, with all models parameterized with the same base LLM and prompts, serving as an ablation study. Response Gen SFT significantly outperforms Response Gen Zero-shot, which suggests that finetuning is required for the LLM to adapt to the response distribution of real student responses. SMART outperforms Response Gen SFT, which validates our hypothesis that preferring responses with high IRT likelihoods in DPO training helps in the difficulty prediction task by calibrating the model in the context of student ability and implicit item difficulty.
@@
Table 1 compares different methods on student simulation metrics. We see that SMART outperforms other methods by a wide margin on the $\theta$-Align metric, which measures the correlation between the desired input ability and inferred ability by scoring the generated responses. This observation shows that DPO training leads to a strong preference for the model to follow the input ability. We also compare the simulated student responses across Response Gen Zero-shot, Response Gen SFT, and SMART, on distribution similarity metrics (MAUVE, FID, and Diversity KL). Response Gen Zero-shot performs poorly, which emphasizes the need for finetuning or DPO training, over simple prompting, to align the LLM with diverse responses by real students. Perhaps surprisingly, Response Gen SFT is marginally better than SMART on these metrics. This observation can be explained by the fact that DPO training prioritizes the IRT likelihoods of responses in its objective; therefore, we get more accurate question difficulty prediction at the expense of text-level similarity, with the model shifting slightly away from the textual distribution to achieve this goal.
@@ 6.2 Qualitative Evaluation
Models Reflect Ability-Specific Features in Student Responses
Tables 2 and 3 show examples of generated responses for students with varying ability levels, for models trained with SFT and with SMART. Broadly speaking, students with low ability have short one-sentence responses, spelling mistakes, poor grammar, and incorrect punctuation. High-ability students, on the other hand, have longer responses with complex sentence structures and strong vocabulary, and make an effort to answer all aspects of the question to receive full credit. SMART also captures student response diversity, especially for medium-ability students, whose responses usually contain simple words, but differ in length, coherence, and score obtained by the trained scoring model. Our preference pair creation encourages this diversity since varied responses can be preferred as long as they are sufficiently likely under the IRT model.
Performance is Item-dependent
We find that SMART’s performance is highly dependent on the type of reading comprehension task (argumentation, narration, explanation, among others) in the item. In particular, difficulty estimates are accurate for items that require information recall or those that add narrative elements to a story. On the other hand, estimates are less accurate for items requiring higher order thinking Lewis and Smith (1993), such as summarizing the main idea of a passage, or writing an introduction to a paragraph that includes a “thesis” or a “main controlling idea”. For these items, we find that SMART tends to overestimate their difficulty. We see that even for high input ability values, SMART tends to generate incomplete responses that only state facts from the passage, rather than summarizing it. Such responses receive only partial credit and thus lead to high difficulty values under IRT. Future work can attempt to augment the training data with more synthetic responses with higher-order thinking, to aid simulation for high-ability students.
Outlier Generated Responses
We observe that SMART occasionally generates responses that are outliers and not found in the training data, with failure patterns including 1) simply repeating the passage, 2) not being relevant to the item, and 3) being highly repetitive or even degenerate. These outlier responses are often inaccurately scored by the scoring model, affecting downstream item difficulty prediction. A possible solution is to use a fidelity checking model to filter out such outlier responses and re-generate when it happens.
@@ 7 Related Work
An increasing body of research uses simulated students to estimate item difficulty. They follow a similar pipeline to our work, where they overgenerate a large set of responses using AI models, and then fit an IRT model to these responses to recover the item difficulty. Such methods have used LSTMs trained with varying proportions of incorrect labels Lalor et al. (2019), prompted ensembles of pre-trained language models to answer the items Liu et al. (2025); Uto et al. (2024), and finetuned language models to predict student responses given prior responses Zelikman et al. (2023). He-Yueya et al. (2024) experiment with LLM ensembles, finetuning, and persona-based prompting, finding that the most effective method depends on the dataset.
Our approach builds on this line of work but introduces several key innovations. First, we focus on open-ended responses, which are not handled by most finetuned or feature-based methods. Second, we explicitly train simulated students to follow a prompted ability level, allowing the simulated students to align with the underlying IRT distribution. Third, our system is designed to operate effectively on small item sets, where feature-based and BERT-based methods typically underperform.
@@ 8 Conclusions and Future Work
There are many avenues for future work. First, we can include generated student responses in the pool of candidates for DPO Parikh et al. (2025), which may help to explicitly disprefer outlier responses that are not aligned with real student data. Second, we can explore the applicability of SMART in other domains such as programming, math, and in non question-response settings such as dialogues Scarlatos et al. (2025). Finally, we can extend SMART’s student representation to go beyond simple scalar-valued student abilities and into more descriptive personas Bui et al. (2025).
@@
There are several practical limitations to our work. First, we only experiment with a single dataset, containing reading comprehension items and student responses at the 6th grade level. While to the best of our knowledge, there are no publicly available open-ended testing datasets like ours, it may be helpful to explore how existing datasets could be adapted to our setting. Second, we restrict our student simulation to models no larger than $3$B parameters, which is relatively small. While our methods are effective at this size, it would be helpful to see if scaling to larger sizes is worth the cost. Third, we find that the textual distribution of our generated responses does not always match the ground-truth responses, and that generated responses are occasionally degenerate. As a result, the pool of simulated responses may not look realistic from a human perspective. This limitation restricts our method’s usefulness as a more generic student simulator, although it could be addressed by improvements to the method in future work.

### v2

@@ 5.3 Baselines
Following prior work AlKhuzaey et al. (2024); Benedetto (2023); Benedetto et al. (2023); Kapoor et al. (2025a); North and Ormerod (2025), we adopt a strong finetuning baseline, Diff SFT, which finetunes a recent and capable encoder-based language model, specifically ModernBERT-base Warner et al. (2024). We use a regression setup, minimizing mean squared error loss, to directly predict item difficulties given the item’s question text and meta data. We also adopt a Random Forest baseline, which is common for difficulty prediction AlKhuzaey et al. (2024); Gombert et al. (2024), where we encode the item’s question text and meta data using the multi-qa-mpnet-base-dot-v1 SBERT model Reimers and Gurevych (2019), and use the resulting embeddings as input features to fit a random forest regressor to item difficulty.
Since our set of items is small, we also introduce several baselines that do not require training. We use a simple yet strong baseline, kNN Mean, that retrieves the $k$ items in the train set with the highest SBERT embedding cosine similarity, and takes their mean difficulty as the predicted test item difficulty. Similar methods have been used in automated scoring Bin et al. (2008); Gomaa and Fahmy (2012). We search over $k\in[1,5]$ and select the best performing value on the validation set, selecting $k=1$ for Smarter Balanced and $k=5$ for CodeWorkout. Finally, we adopt two LLM prompting baselines leveraging in-context learning (ICL) Brown et al. (2020), ICL Random and ICL kNN, where we prompt Llama-3.1-70B-Instruct with $k$ in-context examples of items retrieved from the train set at random or using k-Nearest Neighbors (kNN) Feng et al. (2024), respectively. We ask the model to first reason about the question’s difficulty in a chain-of-thought manner before generating a numeric difficulty prediction. For kNN retrieval, we use the same SBERT model as above, and perform a similar hyperparameter search, selecting $k=4$ for Smarter Balanced and $k=2$ for CodeWorkout.
@@
We compare SMART to other student simulation-based approaches for item difficulty prediction. For a fair comparison, we use the same base model and prompt template as SMART for the following methods, allowing these baselines to also serve as ablations. First, we employ Response Gen ZS, where we prompt an LLM in a zero-shot (ZS) manner to generate a student response given the item and ability. Since the model has not been trained on the distribution of numeric student abilities, we convert abilities to corresponding labels from 10 descriptive buckets shown in Appendix A.2, ranging from “minimal” to “mastery”, with descriptions obtained by prompting Claude 3.5 Sonnet Anthropic (2024). This method is similar to the persona-based prompting method in He-Yueya et al. (2024), where we use ability instead of student demographics. Second, we employ Response Gen SFT, as explained in Section 4.1, which finetunes the LLM on ground-truth student responses given the item and ability. Prior works that finetune simulated students generally include prior student responses in the prompt He-Yueya et al. (2024); Zelikman et al. (2023); we use the student’s ability as a proxy for this information.
Model Smarter Balanced CodeWorkout
PCC $\uparrow$ SCC $\uparrow$ RMSE $\downarrow$ PCC $\uparrow$ SCC $\uparrow$ RMSE $\downarrow$
Direct Item Difficulty Prediction Methods
Diff SFT (ModernBERT) $0.4937$ $0.3188$ $0.7627$ 0.0425-0.0425 0.0036-0.0036 $0.7714$
Random Forest (SBERT) $0.4021$ $0.3115$ $0.8623$ $0.1846$ $0.1418$ $0.6783$
kNN Mean (SBERT) $0.5959$ $0.5219$ $0.6806$ $0.1949$ $0.2415$ $0.6701$
ICL Random (Llama-3.1-70B) $0.1215$ $0.0348$ $1.0426$ $0.2487$ $0.1806$ $0.6462$
ICL kNN (Llama-3.1-70B) $0.5273$ $0.4717$ $0.7634$ $0.0209$ $0.0821$ $0.7397$
Response Gen ZS $0.2651$ $0.2436$ $0.9088$ $0.0733$ $0.0812$ $0.7226$
Response Gen SFT $0.5621$ $0.4254$ $0.7174$ $0.2475$ $0.2388$ $0.6487$
SMART (ours) $\underline{0.6490}$ $0.4861$ $\underline{0.6699}$ $0.3082$ $0.3188$ $0.6225$
Response Gen ZS $0.3446$ $0.3963$ $0.8568$ $0.1487$ $0.1636$ $0.6912$
Response Gen SFT $0.6069$ 0.5830 $0.6905$ 0.4256 $\underline{0.3648}$ 0.5627
SMART (ours) 0.6737 $\underline{0.5661}$ 0.6197 $\underline{0.3929}$ 0.4230 $\underline{0.5890}$
Table 1: Performance on question difficulty prediction. Best performance is in bold and second best is underlined. SMART outperforms all direct prediction methods and generally outperforms other simulation-based methods.
@@ 6 Results, Analysis, and Discussion
We now detail our quantitative results on item difficulty prediction and student response simulation, and qualitatively analyze the simulated student responses and SMART’s strengths and weaknesses.
@@ 6.1 Quantitative Results
We report the average performance across all test items and cross-validation folds for both datasets in Table 1. Our approach, SMART, outperforms all direct prediction methods on all metrics, even largely maintaining this advantage when using the small 1B base LLM. It also outperforms all other student simulation methods, with the exception of SFT when using 3B models; in these cases, SMART outperforms SFT by wide margins on some metrics while it is closely behind on others. These results indicate that both methods are powerful predictors of item difficulty, although there is high variability due to the small number of items in each dataset; larger test sets may be required to more clearly show SMART’s advantage. We also observe that CodeWorkout is a more challenging dataset, with most methods performing worse than they do on Smarter Balance in terms of PCC and SCC. Still, the fact that SMART consistently outperforms baselines on both datasets shows that it can generalize across educational domains and student populations.
Model Smarter Balanced CodeWorkout
MAUVE $\uparrow$ FID $\downarrow$ Div. KL $\downarrow$ $\theta$-Align $\uparrow$ MAUVE $\uparrow$ FID $\downarrow$ Div. KL $\downarrow$ $\theta$-Align $\uparrow$
Response Gen ZS $0.0058$ $0.2713$ $0.2535$ $0.0399$ $0.0196$ $0.0747$ $0.0714$ 0.0014-0.0014
Response Gen SFT $0.1483$ $0.0847$ $\underline{0.0215}$ $0.4481$ $\underline{0.1777}$ $\underline{0.0269}$ $\underline{0.0315}$ $0.1730$
SMART (ours) $0.1315$ $0.0918$ $0.0482$ $\underline{0.7338}$ 0.1931 0.0267 0.0284 0.5895
Response Gen ZS $0.0052$ $0.3386$ $0.4032$ $0.0211$ $0.0042$ $0.1284$ $0.2779$ 0.0101-0.0101
Response Gen SFT 0.1858 0.0757 0.0137 $0.5158$ $0.1426$ $0.0333$ $0.0515$ $0.2226$
SMART (ours) $\underline{0.1713}$ $\underline{0.0785}$ $0.0279$ 0.7655 $0.1479$ $0.0303$ $0.0335$ $\underline{0.5582}$
Table 2: Performance on student simulation metrics. Best performance is in bold and second best is underlined. SMART significantly outperforms other methods on ability following.
@@
We observe that direct prediction baselines significantly underperform student simulation methods on both datasets. Diff SFT, which follows the approach of existing state-of-the-art methods AlKhuzaey et al. (2024); Benedetto (2023); Benedetto et al. (2023), performs moderately well on Smarter Balanced but poorly on CodeWorkout. This result is perhaps not surprising, since Diff SFT’s training signal completely comes from given item difficulties, and thus can only work when a large number of training items are available. Direct difficulty prediction baselines that require no training also follow a similar pattern: kNN Mean and ICL kNN perform moderately on Smarter Balanced, likely because reading comprehension questions with similar semantic features are likely to pose similar tasks to the student and thus have similar difficulties. However, since this heuristic does not apply to coding questions, these baselines do not generalize across datasets. The poor performance of ICL Random also shows that powerful LLMs do not perform well on the difficulty prediction task without careful calibration. In our low resource setting with few items to train with, student simulation is key: student ability and item difficulty are only related via the response score when training an IRT model, but our student simulator can effectively couple IRT training with textual student responses and a scoring model. As a result, this extra information helps us to calibrate the IRT model better, making it easier to generalize to previously unseen items.
@@
We compare the performance of Response Gen ZS, Response Gen SFT, and SMART, with all models parameterized with the same base LLM and prompts, serving as an ablation study. Response Gen SFT significantly outperforms Response Gen ZS, which suggests that finetuning is required for the LLM to adapt to the response distribution of real student responses. With the exception of some metrics when using 3B models, SMART outperforms Response Gen SFT, which validates our hypothesis that preferring responses with high IRT likelihoods in DPO training helps in the difficulty prediction task by calibrating the model in the context of student ability and implicit item difficulty.
@@
Table 2 compares the student simulation methods on metrics regarding the quality and faithfulness of simulated responses. We see that SMART outperforms other methods by a wide margin on the $\theta$-Align metric, which measures the correlation between the desired input ability and inferred ability by scoring the generated responses. This observation shows that DPO training leads to a strong preference for the model to follow the input ability, resulting in a faithful student simulator. We also compare the simulated student methods on distribution similarity metrics, measuring how closely their generated responses align with ground-truth ones. Response Gen ZS performs poorly, which emphasizes the need for finetuning or DPO training, over simple prompting, to align the LLM with diverse responses by real students. Interestingly, Response Gen SFT is marginally better than SMART on these metrics on Smarter Balance, while SMART is better on CodeWorkout. This observation highlights the fact that DPO training prioritizes the IRT likelihoods, which may or may not align with how realistic the generated responses appear. We also observe that 1B models outperform 3B models on these metrics on CodeWorkout, perhaps due to the lower coding ability of the 1B model, therefore aligning better with code written by novice coders (students).
@@ 6.2 Qualitative Evaluation
Models Reflect Ability-Specific Features in Student Responses.
Tables 3 and 4 show examples of generated responses to Smarter Balanced items for students with varying ability levels. Compared to SFT, SMART more clearly reflects realistic patterns based on ability. Broadly speaking, students with low ability have short one-sentence responses, spelling mistakes, poor grammar, and incorrect punctuation. High-ability students, on the other hand, have longer responses with complex sentence structures and strong vocabulary, and make an effort to answer all aspects of the question to receive full credit. SMART also captures student response diversity, especially for medium-ability students, whose responses usually contain simple words, but differ in length, coherence, and score obtained by the trained scoring model. Our preference pair creation encourages this diversity since varied responses can be preferred as long as they are sufficiently likely under the IRT model.
Table LABEL:tab:qual_ex_codeworkout shows examples of real and generated responses to a CodeWorkout item. SMART’s difficulty prediction is much more accurate than SFT’s on this item, primarily due to the diversity of generated responses; the SFT responses are semantically similar and receive similar scores across ability levels, while for SMART, low ability students score poorly while medium and high ability students score highly. Additionally, SMART more faithfully reflects semantic patterns in the student code, such as medium ability responses being longer and high ability responses being more concise and using more clear coding constructs.
Performance is Item-dependent.
For Smarter Balanced items, we find that SMART’s performance is highly dependent on the type of reading comprehension task (argumentation, narration, explanation, among others) in the item. In particular, difficulty estimates are accurate for items that require information recall or those that add narrative elements to a story. On the other hand, estimates are less accurate for items requiring higher order thinking Lewis and Smith (1993), such as summarizing the main idea of a passage, or writing an introduction to a paragraph that includes a “thesis” or a “main controlling idea”. For these items, we find that SMART tends to overestimate their difficulty. We see that even for high input ability values, SMART tends to generate incomplete responses that only state facts from the passage, rather than summarizing it. Such responses receive only partial credit and thus lead to high difficulty values under IRT. Future work can attempt to augment the training data with synthetic responses with higher-order thinking to aid simulation for high-ability students.
Outlier Generated Responses.
We observe that SMART occasionally generates outlier responses with fidelity issues, with failure patterns including 1) repeating the passage, 2) not being relevant to the item, and 3) being highly repetitive. These outlier responses are often inaccurately scored by the scoring model, affecting downstream item difficulty prediction. A possible solution is to use a fidelity checking model to filter out such outlier responses and re-generate when it happens.
@@ 7 Related Work
An increasing body of research uses simulated students to estimate item difficulty. They follow a similar pipeline to our work, where they overgenerate a large set of responses using AI models, and then fit an IRT model to these responses to recover the item difficulty. Such methods have prompted ensembles of pre-trained language models to answer the items Liu et al. (2025); Uto et al. (2024), finetuned language models to predict student responses given prior responses Zelikman et al. (2023), and trained LSTMs with varying proportions of incorrect labels Lalor et al. (2019). He-Yueya et al. (2024) experiment with LLM ensembles, finetuning, and persona-based prompting, finding that the most effective method depends on the dataset.
Our approach builds on this line of work but introduces several key innovations. First, we explicitly train simulated students to follow a prompted ability level, allowing the simulated students to align with the underlying IRT distribution. Second, we focus on open-ended responses, which are not handled by most finetuned or feature-based methods. Third, our system is designed to operate effectively on small item sets, where feature-based and BERT-based methods typically underperform.
@@ 8 Conclusions and Future Work
There are many avenues for future work. First, we can include generated student responses in the pool of candidates for DPO Parikh et al. (2025), which may help to explicitly disprefer outlier responses that are not aligned with real student data. Second, we can explore the applicability of SMART in other domains such as programming, math, and in non question-response settings such as dialogues Scarlatos et al. (2025b). Finally, we can extend SMART’s student representation to go beyond simple scalar-valued student abilities and into more descriptive personas Wu et al. (2025).
Acknowledgments
We thank Smarter Balanced for providing the private dataset used in our experiments. We also thank Joseph Di Garbo, Frank Rijmen, Suhwa Han, Benjamin Godek, Bokhee Yoon, Zhangqi Duan, and Jacqueline Scarlatos for helpful discussions around this work. This work is partially funded by the NSF under grants 2237676 and 2418657.
@@ Acknowledgments
There are several practical limitations to our work. First, while we experiment on two datasets across educational domains and student populations, there are limitations to these datasets. The Smarter Balanced dataset is not publicly available, and the CodeWorkout data takes place over the course of two semesters, making static ability modeling challenging. Additionally, the small number of items in each dataset makes reliable evaluation challenging. To the best of our knowledge, there are no other suitable publicly available open-ended testing datasets. Second, we restrict our student simulation to models no larger than 33B parameters, which is relatively small. While our methods are effective at this size, it would be helpful to see if scaling to larger sizes is worth the cost. Third, we find that the textual distribution of our generated responses does not always match the ground-truth responses, and that generated responses are occasionally degenerate. As a result, the pool of simulated responses may not look realistic from a human perspective. This limitation restricts our method’s usefulness as a more generic student simulator, although it could be addressed by improvements to the method in future work.